OpenStudy (highschoolmom2010):

A fire drill will begin at a randomly-chosen time between 8:30 a.m. and 3:30 p.m. You have a math test planned for 2:05 p.m. to 3:00 p.m. If the fire drill is in the afternoon, what is the probability that it will start during the test?

ganeshie8 (ganeshie8):

Event A : firedrill to start in afternoon Event B : firedrill to start during test

ganeshie8 (ganeshie8):

P(B | A) = 25minutes /(3*60+30)minutes = 25/210 = 5/42

OpenStudy (highschoolmom2010):

any ideas how my book got 11/42

ganeshie8 (ganeshie8):

okay, ur textbook is correct again. i made a mistake

ganeshie8 (ganeshie8):

mistake in counting # of minutes the test takes :/

ganeshie8 (ganeshie8):

how many minutes are there between : 2:05 p.m. to 3:00 p.m ?

OpenStudy (highschoolmom2010):

55

ganeshie8 (ganeshie8):

yes, and how many minutes are there between "afternoon" and "3:30" ?

ganeshie8 (ganeshie8):

btw, noon = 12:00

ganeshie8 (ganeshie8):

12 to 3:30 that 3 and half hours right ?

OpenStudy (highschoolmom2010):

210

ganeshie8 (ganeshie8):

yes, so probability for the drill to occur during test, given that drill occured after noon :- P(B | A) = 55/210

ganeshie8 (ganeshie8):

simplify

OpenStudy (highschoolmom2010):

oh ok that makes more sense now

OpenStudy (highschoolmom2010):

ty @ganeshie8

ganeshie8 (ganeshie8):

np :)